In 4 years, Mike will be 3 times as old as Chris will be. The sum of their ages now is 56. I just need to know how to set this equation up! Is there a method for all problems like this?
Answer by rwm(914)

(m+4)=3*(c+4)
m+c=56
c=12 m=44
in 4 years
check
44+4=48=3*(12+4)=3*16=48
ok